Please help me solve this question:Give the next three terms of the sequence -8,-5,-2,1?
Answer by Theo(13342)

add 3 to -8 and you get -5.
add 3 to -5 and you get -2.
add 3 to -2 and you get 1.

it appears you have an arithmetic progression where the common difference is 3.

if you add 3 to each succeeding term, you'll have your solution.

1 + 3 = 4
4 + 3 = 7
7 + 3 = 10

your entire sequence looks like this:

-8 + 3 = -5
-5 + 3 = -2
-2 + 3 = 1
1 + 3 = 4
4 + 3 = 7
7 + 3 = 10